[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Oct 31, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Nov 1 03:55:42 UTC 2016


HSU Shift Log for Mon Oct 31 2016 23:55:36 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None. However, TAFB received call from Tech Control  
(NCO?) around 600 pm about our ftp line being down; however, access to NHC  
ftp link was working fine; TAFB contacted TSB for clarification.

Miscellaneous issues: NHC web page updated NHC TWOAT/EP and TAFB TWD  
graphics correctly at 800 pm. System appears to be functioning properly.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161101/e707afc9/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list